Session: Feb 1999 (session 1)
Wavelength: 18 cm


To give feedback on this experiment, click here

Experiment Feedback

Effelsberg (dgraham at Wed Feb 24 09:45:46 MET 1999) Minor failures
LCP RCP slightly different noise since bandedge. Stopped due to snow from 1314-1351,1411-1444,1649-1710,2138-2146.

Onsala (A. Polatidis at Wed Feb 24 13:27:34 MET 1999) Minor failures
Lost data between 2117 and 2152 because the first tape got mangled in the vacuum chamber during the tape change. There was no damage to the first tape, only the first couple of feet. No other problems.

Westerbork (Tony Foley at Wed Feb 24 19:52:23 MET 1999) Minor failures
10:00 to 11:30 UT lost because of bad procedures. 18:16 to 19:03 lost because the tape lost vacuum (lots of dirt off tape blocked vacuum sensors). Very many scans out of HA range or below horizon.

Calibration (Jean\-Francois Desmurs at Wed Mar 3 11:25:44 MET 1999) Success
The ANTAB formatted file for this experiment can be found on VLBEER in the feb99 directory under the name: gp021.antab. Read the comments in the file to get further information.